OpenStudy (anonymous):

The equation of a line is 2x - 2y = 6. What is the slope of the line? A. -3 B. -1 C. 1/3 D. 1

OpenStudy (hanner_b_nanner):

the slope will be 1!! :))

OpenStudy (anonymous):

How would you figure that out?

OpenStudy (anonymous):

I am doing a study guide for my test and I need to make sure I know.

OpenStudy (hanner_b_nanner):

I just did it the easy way, and graphed it! It comes out to rise one, run one.. and it's going up from the negative, towards the positive! You could just change it to slope intercept form to find the slope, but doing it this way is so much easier! :))

OpenStudy (anonymous):

Oh okay! Thanks :)

OpenStudy (hanner_b_nanner):

anytimeee!!

OpenStudy (popitree):

slope form: y = mx+c and m is slope, here 2y=2x-6 or y = 1.x - 3 so slope= 1

OpenStudy (hanner_b_nanner):

yes^^ thanksss!! :D